Your goal is to have $17,500 in your bank account by the end of eight years. If the interest rate remains constant at 9% and you want to make annual identical deposits, how much will you need to deposit in your account at the end of each year to reach your goal? (Note: Round your answer for PMT to two decimal places.)

If your deposits were made at the beginning of each year rather than an at the end, by how much would the amount of your deposit change if you still wanted to reach your goal by the end of eight years? (Note: Round your answer for PMT to two decimal places.)
Answer #1

This question requires application of FV of ordinary annuity and annuity due (where payments are made at the beginning of year).

For first part, where it is an ordinary annuity, formula would be,

[(1+r) -1 FV of Annuity = P P= Periodic Payment r=rate per period n = number of periods

17,500 = P* (1 +0.09) - 1 0,09

17,500 = P * 11.0285

P = $1,586.80 --> Answer to part 1

For second part, where it is an annuity due, formula would be,

- 1] FV of Annuity Due = (1 + r) x P P= Periodic Payment r=rate per period n = number of periods

- 1 17,500 = (1+0.09) *P* (1 + 0.09) 0.09

17,500 = 1.09 * P * 11.0285

P = $1,455.78

Change = $1,586.80 - $1,455.78 = $131.02 ---> Answer to Part 2